OpenStudy (anonymous):

Did i solve this problem correctly? Translate the following statement into an open sentence using absolute value and the variable x:? The numbers whose distance from -2 is at least 5 units. |x-2|=5

OpenStudy (accessdenied):

When I test with -3 (which is 1 unit from -2), I get that |-3-2| = |-5| = 5 units. 1) Since its from -2, you should take x - (-2), or x + 2 instead. That way, |-3 + 2| = |-1| = 1 2) It says "at least 5 units", so you can have more than 5 as well. The equality should be |x+2| >= 5

OpenStudy (accessdenied):

\( >= \) being \( \ge \), that might make it easier to read. *